find the values of x and y that solve the following system of equations
-3x- 8y=19
4x+9y=-22

Answer by mananth(16946)

-3x- 8y=19.................1
4x+9y=-22...................2
..
multiply 1 by 4 & 2 by 3
..
-12x-32y=76
12x+27y=-66
add them
-32y+27y =76-66
-5y=10
/-5
y = -2
...
plug value of y in 1
-3x-8*-2=19
-3x+16=19
-3x=19-16
-3x=3
/-3
x =-1
...
m.ananth@hotmail.ca